SIPB Installer for Athena Linux Coming Soon

August 29, 2000

Linux is a freely distributable implementation of UNIX. At MIT, our public computer clusters run a standard set of software known as Athena. You may have noticed Linux-Athena workstations in the public clusters. These are Intel X86-based workstations running a port of Athena to the Red Hat distribution of Linux that is maintained by Information Systems. The software running on those workstations is available to the MIT community, but it is currently difficult to install it on hardware which is not identical to that which is in the clusters.

SIPB is working to modify the installer shipped with Red Hat Linux to make installing a Linux-based Athena system on your personal computer relatively straightforward. We had hoped to have this installer ready in time for the start of the academic year, but unfortunately it is still in testing. It should be ready in a few weeks.

If you would like to be notified when the installer is ready, subscribe to the mailing list linux-announce by running, at the athena% prompt,

        blanche linux-announce -a $USER

This list is also used for important announcements related to older releases of Athena for Red Hat Linux that were entirely maintained by SIPB. Those releases are based on older versions of Red Hat and Athena software so we recommend that you wait for our new installer.

If you are buying a computer, and are thinking of running Linux on it, you can buy a machine just like the cluster ones, or you can look at the list of PCs that MIT recommends for Windows users. See

        http://web.mit.edu/is/desktop/saws_configs.html

for the cluster configuration, and

        http://web.mit.edu/is/desktop/pcadmin.html

for the MIT recommended hardware. Any of the listed configurations will work fine with our forthcoming installer.

If you already have a computer and only need to purchase an Ethernet card to connect it to MIT’s network, please see

        http://web.mit.edu/is/desktop/pcether.html

If you’d like to learn more about Red Hat Linux, a useful web site is

        http://www.redhat.com/support/docs/rhl62.html

It contains a frequently asked questions list and an introduction to Linux. Note that you should not install Linux on your computer until our installer is released. For more information

about Athena, please see http://web.mit.edu/olh/.

Meanwhile, we’d still like to hear from you. Feel free to stop by our office (W20-557) and ask any questions you might have, or email sipb@mit.edu.